OpenStudy (anonymous):

Which of the following problem types can always be solved using the law of cosines or sines? A. SSS B. SSA C. SAS D. SAA E. AAA F. ASA choose all that apply

OpenStudy (anonymous):

YMCA

OpenStudy (anonymous):

Law of sines is: Sin(A)/a = Sin(B)/b You need three of those variables to solve for the fourth, so you need 2 angles and a side. Since you have 2 angles, you also know the third, which allows you to solve for the 3rd side. Law of cosines is: c^2 = a^2 + b^2 - 2ab*cosC There are 4 variables. 3 sides and an angle. If you have 3 sides, you can solve for the angles. If you have 2 sides and an angle opposite one, you can solve for the 3rd side.